The ratio of a two-digit number and the number and the number obtained by interchanging the digits is 4:7. If the difference of the digits is 3. Find the number. (Ans. 36)
Answer by lwsshak3(11628)

The ratio of a two-digit number and the number and the number obtained by interchanging the digits is 4:7. If the difference of the digits is 3. Find the number. (Ans. 36)
**
let u=units digit
let t=ten's digit
original number=10t+u
interchanged number=10u+t
u-t=3
u=3+t
..
(10t+u)/(10u+t)=4/7
70t+7u=40u+4t
66t=33u=33(3+t)=99+33t
33t=99
t=3
u=3+t=6
ans:
number=tu=36